# The LUX-ZEPLIN (LZ) Collaboration Study of few-electron backgrounds in the LUX-ZEPLIN detector

### v1

@@ I Introduction
To detect dark matter particles with masses below a few GeV/, an experiment must be sensitive to interactions depositing energies at the keV-level and below [5]. In this energy region, S1 signals from both electron and nuclear recoils are too weak to be observed due to the relatively low light detection efficiency in TPCs (typically around 10%). However, S2 signals can be collected with high enough efficiency that the extraction of single ionization electron signals may be detected. This allows the energy threshold of a detector to be lowered for ionization-only dark matter searches [6, 7, 8], compared to searches requiring both S1 and S2 signals. When combined with the hypothesized Migdal effect [9, 10], ionization-only searches using LXe TPC experiments including XENON1T, LUX and PandaX have probed dark matter masses as low as a few MeV [11, 12, 13].
@@
Spontaneous emission of single and multi-electron (ME) pulses from high voltage grid wires is a well-documented effect [33, 34, 35]. The most commonly considered sources for this background are wire defects and debris on grid wires. To reduce the chances of debris being deposited on grid wires, extensive cleaning and dust control protocols were implemented during the fabrication and handling of LZ grids [34]. To counteract the presence of microscopic wire defects, the LZ TPC extraction (gate) grid was passivated with citric acid after fabrication to remove a thin surface layer and form a homogeneous oxide layer. Despite these precautions, intense electron emission from “hot spots” was occasionally observed during TPC operation even with otherwise stable high voltage conditions. Fowler-Nordheim (F-N) emission [36] has been proposed as a possible mechanism, but observed emission behaviors are not consistent with F-N theory [35]. Namely, emission rates do not monotonically increase with grid voltages, and the electric field on the wire surfaces is not expected to reach the level required for significant F-N emission even with enhancements from defects [34]. A hot spot analysis cut was used by LZ to reject data acquired when one or more significant emission centers were active [3]. This work characterizes the electron emission behavior from these hot spots and also investigates new methods to tag emission pulses so these backgrounds may be rejected. High-voltage grids can also emit electrons as a result of radioactive decays on the grid wire surfaces, which may produce S2s containing up to thousands of electrons [37]; this phenomenon is not a focus of this work.
@@ III Delayed electron emission
After each progenitor event, all small electron pulses are identified in subsequent data acquisition windows, and the process stops when any pulse larger than 100 (“vetoing pulse”) is detected. Progenitors followed by a vetoing pulse within 5 ms are discarded. We also require a progenitor to be at least 200 ms after any vetoing pulse to mitigate the aftermath of earlier energy depositions. Finally, events with otherwise valid progenitors are rejected if any data buffer for storing PMT digitizations is saturated. This cut effectively excludes progenitor pulses larger than phd and largely mitigates the effects of PMT saturation.
@@
After 3 ms, the pulses are almost exclusively SEs, the rate of which is more than an order of magnitude higher than that of pulses tagged as double electrons. The rate of pulses identified as ME ( 3 electrons) are statistically consistent for the position-correlated and position-uncorrelated distributions. This suggests that the dominant source of the ME background is independent of preceding energy depositions. The position-correlated SE rate, shown in Fig. 2, greatly exceeds that of position-uncorrelated SEs, so we attribute the position-correlated SEs to delayed emission related to earlier progenitor events. Similar to prior studies, the position-correlated SE rate approximately follows a power-law until approaches 200 ms, after which contributions from position-uncorrelated pulses become significant. Therefore, we subtract the rate of position-uncorrelated pulses from the rate of position-correlated pulses in each bin, and fit the resulting distribution with the function , where and are constants, for times between 3 and 200 ms. At the nominal WS2022 field values we obtain a value of with negligible changes in value from varying the choice of fit bounds between 3–200 ms. When analyzing datasets acquired under the same nominal conditions but at times several months apart, we observe a 10% variation in , which we attribute to subtle changes in detector conditions and treat as a systematic error for this analysis. The obtained value and uncertainty level are in agreement with the values reported in LUX [25] and XENON1T [27], and will be discussed further in the next subsections.
@@ III.1 Impurity and drift time dependence
| Factor | Description |
| = | Raw S2 area divided by raw SE area |
| = | Corrected for extraction efficiency |
| = | Corrected for drift electron loss |
| = | Drift electron loss |
@@ III.2 Electric field dependence
Table 2 lists all prior reports of the electron train power law in dual-phase TPCs. An exponent value ranging from -1 to -1.4 is reported across different experiments. Consistent with our analysis shown in Fig. 3, the liquid xenon purity, measured as the electron lifetime, does not seem to have significant bearing on the power law slope. In addition, these measurements were carried out with varying electron extraction efficiency values, which also do not seem to impact the power-law exponent. In LZ, we analyzed data acquired at a fixed drift field (193 V/cm) but three different extraction fields: 3410, 3900, and 4300 V/cm, corresponding to extraction efficiencies of 50–78%. The power-law exponent values measured at these different field settings were consistent with each other.
@@
Due to the short drift distance for progenitor ionization signals, very few electrons are captured by impurities in this volume, resulting in low statistics of delayed electrons. However, after subtracting the position-uncorrelated electron rate, we observe clear power-law features up to a 20 ms delay time window. We fit power laws to the background-subtracted rates at both field settings with a fit range restricted to 3–20 ms in order to avoid contamination from the photoionization cascade. At 3.4 kV/cm, we obtain and at 3.9 kV/cm we obtain . These results, along with the others reported in this work, are summarized in Tab. 2.
@@
Until now, there has been no compelling hypothesis for the mechanism behind delayed electron emission. The non-integer power-law exponent observed in LZ and other experiments makes it difficult to attribute this phenomenon to processes such as diffusion [44] or recombination [45], which can lead to or time dependence. Non-integer power laws are nevertheless observed in systems with energy traps of varying depths [46]. Taking the example of pumping out a vacuum chamber after it is exposed to air [47, 48], when there is only one residual gas species with a specific surface adsorption energy, the chamber pressure will experience an exponential decay when it is pumped on. However, when a range of adsorption energy values are introduced, the overall gas pressure exhibits a power-law time dependence. In liquid xenon, Ref. [32] proposes that thermal collisions with neighboring atoms may cause negatively charged impurities to be re-ionized and electrons to be released. In principle, these interactions should also result in an exponential time dependence for the electron emission rate, but the presence of multiple species of impurities with varying electron affinity can lead to an overall emission rate in the form of a power-law of non-integer exponents. Furthermore, the thermal energy of charged particles in an electric field depends on the field strength, so the hypothesis that charged impurities release electrons as a result of thermal collision can be tested by a definitive field-dependence study in the future.
@@ IV.1 Photons accompanying grid electron emission
This section studies electron emission from the LZ gate electrode using an extraction region-only dataset acquired during commissioning. This configuration greatly reduces delayed electron backgrounds following particle interactions in the drift region and is therefore ideal for characterizing grid emission. To further reduce the delayed electron background from interactions in the above-gate liquid, we veto time periods of following large S2s, where the veto length increases with the size of the S2 [3]. We also require that the S2 pulses do not have S1s larger than 100 phd within 3 s beforehand to exclude photoionization electrons. After applying this selection and additional position reconstruction quality cuts, the X-Y position distribution of the remaining isolated electron pulses is shown in Fig. 6 (top). Localization of the pulses around specific “hot spots,” or spontaneous emitters, is clearly visible. In addition to the spatial inhomogeneity, large temporal spikes in the S2 rates are also present in this data set, as shown in Fig. 6 (bottom). Because the cathode is inactive, the hot spots are attributed to electron emission from defects or debris on the gate electrode. During the enhanced electron emission periods, high rates of photons are also observed, a phenomenon also reported in LUX [35]. Figure 7 (top) shows the arrival time of photons relative to that of S2s, calculated as the number of single photo-electrons (SPEs) per S2. SPEs are defined in the LZ reconstruction algorithm as S1-like signals with light in only one PMT. A prominent peak is observed at 2.2 s, which is consistent with the drift time from the gate wire to the liquid surface at the WS2022 field. As described in Sec. III.2, the maximum drift time for extraction region events is approximately 2.5 s, but the drift time for gate-emitted electrons also depends on the exact emission position on the wire and the local field structure. This observation confirms that these photons and electrons are emitted around the same time and that the electrons originate from the gate grid. Similar gate photon-electron coincidences are also observed in a study of radiogenic backgrounds on LZ grids [37]; the rate of grid radioactive decays, however, is orders of magnitude lower than that observed for field-induced emission in the few-electron region and thus its contribution is neglected.
@@ IV.2 Implications for dark matter searches
The vast majority of grid-emitted S2 pulses observed in LZ (Fig. 8) contain less than 5 electrons (300 phd). The standard WIMP search of LZ uses an S2 threshold of approximately 10 electrons [3], so grid emission only contaminates the region of interest when rates are high enough to pile up significantly. In WS2022 [3], we implemented a rate-based hot spot exclusion cut by removing time intervals (120 second) where the average rate of low-energy S2s (including SEs) exceeded a certain threshold. This cut was demonstrated to be sufficient in reducing spurious S2 backgrounds for a sensitive search of medium to high-mass dark matter candidates.
@@
For low-mass dark matter searches, we usually focus on events with no identified S1s111The S1 definition is the same as that used in WS2022 where at least 3 photomultiplier tubes must observe light. within the 4.5 ms event window, and this cut is found to accept of S2-only events between 10 and 500 electrons. The signal loss is dominated by false SPE tagging, which we further investigate with a different method. The WS2022 data, with the TPC drift region active, contains a higher SPE rate than that in the extraction region-only dataset shown in Fig. 7. As a result, the probability for a SPE to be randomly paired with a S2 pulse can be significant and cause non-grid S2 pulses to be falsely rejected. We carried out a mock analysis by tagging SPEs arriving in a control drift time window that is equal in size but offset from the gate cut interval. Then we remove all S2 pulses in an event if any S2 pulse of 300 phd has a tagged SPE in the control coincidence window. The SPE-S2 pairs tagged in this way are only from random coincidence and the resulting event rate reduction is a measure of signal loss from random photon coincidence. When this mock cut is applied to S2-only events, we found a survival rate of for events with S2 pulses above 300 phd (and below 2,880 phd, or 50 electrons), and a lower rate of for events with S2 pulses below 300 phd. The survival rates are different for S2 pulses above and below 300 phd because a large S2 can only be rejected in this cut if it shares the event window with a tagged S2 of 300 phd, but a small S2 can be additionally rejected if it has a preceding SPE pulse.
The uncertainties quoted above are statistical only, and a full acceptance analysis is not attempted here because it would depend on data trigger efficiency and complex timing structures of SPE and S2 rates. The obtained efficiency values, nevertheless, suggest that a substantial reduction in grid backgrounds can be achieved with a a high signal acceptance of 90% or more. It is worth noting that this proof-of-principle analysis does not use the standard LZ hot spot cut that excludes events recorded during periods of high average S2 rates. There is considerable overlap between the background populations removed by these two cuts. However, the event-level method developed in this work can enhance sensitivity to weak grid-emission with an average rate too low to be detected by rate-based cuts. This cut can be further optimized to balance background rejection power and signal loss by tuning the width of pulse rejection window and the minimum number of photon-tagged S2s within the window. One can further exploit the position distribution of tagged S2s to further reduce accidental coincidences and to improve the signal efficiency. The optimization of this background rejection method can be explored in future work.
@@ Acknowledgements.
- Akerib et al. [2022] D. S. Akerib, P. B. Cushman, C. E. Dahl, R. Ebadi, A. Fan, R. J. Gaitskell, C. Galbiati, G. K. Giovanetti, G. B. Gelmini, L. Grandi, S. J. Haselschwardt, C. M. Jackson, R. F. Lang, B. Loer, D. Loomba, M. C. Marshall, A. F. Mills, C. A. J. OHare, C. Savarese, J. Schueler, M. Szydagis, V. Takhistov, T. M. P. Tait, Y. D. Tsai, S. E. Vahsen, R. L. Walsworth, and S. Westerdale, “Snowmass2021 cosmic frontier dark matter direct detection to the neutrino fog,” (2022), arXiv:2203.08084 [hep-ex].
@@
- Essig et al. [2023] R. Essig et al., “Snowmass2021 cosmic frontier: The landscape of low-threshold dark matter direct detection in the next decade,” (2023), arXiv:2203.08297 [hep-ph].
@@
- Li et al. [2023] S. Li et al. (PandaX Collaboration), Physical Review Letters 130 (2023), 10.1103/physrevlett.130.261001.
@@
- Ibe et al. [2018] M. Ibe, W. Nakano, Y. Shoji, and K. Suzuki, JHEP 03, 194 (2018), arXiv:1707.07258 [hep-ph].
@@
- Aprile et al. [2022a] E. Aprile et al. (XENONnT Collaboration), The European Physical Journal C 82 (2022a), 10.1140/epjc/s10052-022-10345-6.
@@
- Kopec et al. [2021] A. Kopec, A. Baxter, M. Clark, R. Lang, S. Li, J. Qin, and R. Singh, Journal of Instrumentation 16, P07014 (2021).
- Aprile et al. [2022b] E. Aprile et al. (XENON Collaboration), Physical Review D 106 (2022b), 10.1103/physrevd.106.022001.
- Akimov et al. [2016] D. Akimov, V. Belov, A. Bolozdynya, A. Burenkov, Y. Efremenko, A. Etenko, V. Kaplin, A. Karelin, A. Khromov, M. Kirsanov, S. Klimanov, A. Kobyakin, A. Konovalov, A. Kovalenko, A. Kuchenkov, A. Kumpan, Y. Melikyan, R. Nikolaev, D. Rudik, G. Simakov, V. Sosnovtsev, and V. Stekhanov, Journal of Instrumentation 11, C03007 (2016).
@@
- Sorensen and Kamdin [2018] P. Sorensen and K. Kamdin, Journal of Instrumentation 13, P02032 (2018), arXiv:1711.07025 [astro-ph, physics:hep-ex, physics:physics].
@@
- Mount et al. [2017] B. J. Mount et al. (LZ Collaboration), (2017), arXiv:1703.09144 [physics.ins-det].
@@
- Chiggiato [2020] P. Chiggiato, “Outgassing properties of vacuum materials for particle accelerators,” (2020), arXiv:2006.07124 [physics.acc-ph].

### v2

# Study of few-electron backgrounds in the LUX-ZEPLIN detector

@@ I Introduction
To detect dark matter particles with masses below a few GeV/, an experiment must be sensitive to interactions depositing energies at the keV-level and below [5]. In this energy region, S1 signals from both electron and nuclear recoils are too weak to be observed due to the relatively low light detection efficiency in TPCs (typically around 10%). However, S2 signals can be collected with high enough efficiency that the extraction of single ionization electron signals may be detected. This allows the energy threshold of a detector to be lowered for ionization-only dark matter searches [6, 7, 8, 9, 10], compared to searches requiring both S1 and S2 signals. When combined with the hypothesized Migdal effect [11, 12], ionization-only searches using LXe TPC experiments including XENON1T, LUX and PandaX have probed dark matter masses as low as a few MeV [13, 14, 15].
@@
Spontaneous emission of single and multi-electron (ME) pulses from high voltage grid wires is a well-documented effect [35, 36, 37]. The most commonly considered sources for this background are wire defects and debris on grid wires. To reduce the chances of debris being deposited on grid wires, extensive cleaning and dust control protocols were implemented during the fabrication and handling of LZ grids [36]. To counteract the presence of microscopic wire defects, the LZ TPC extraction (gate) grid was passivated with citric acid after fabrication to remove a thin surface layer and form a homogeneous oxide layer. Despite these precautions, intense electron emission from “hot spots” was occasionally observed during TPC operation even with otherwise stable high voltage conditions. Fowler-Nordheim (F-N) emission [38] has been proposed as a possible mechanism, but observed emission behaviors are not consistent with F-N theory [37]. Namely, emission rates do not monotonically increase with grid voltages, and the electric field on the wire surfaces is not expected to reach the level required for significant F-N emission even with enhancements from defects [36]. A hot spot analysis cut was used by LZ to reject data acquired when one or more significant emission centers were active [3]. This work characterizes the electron emission behavior from these hot spots and also investigates new methods to tag emission pulses so these backgrounds may be rejected. High voltage grids can also emit electrons as a result of radioactive decays on the grid wire surfaces, which produce a continuous S2 spectrum up to thousands of electrons [39, 40]. However, the low collection efficiency for electrons on the cathode grid surfaces, where most field lines divert away from the drift region, can cause high energy events to appear in the small-S2 region. Additionally, analysis cuts to reject non-bulk events often lose power for small S2s, and as a result, radiogenic grid backgrounds have been observed to contribute significantly to the small-S2 region [14, 29, 10].
@@ III Delayed electron emission
After each progenitor event, all small electron pulses are identified in subsequent data acquisition windows, and the process stops when any pulse larger than 100 (“vetoing pulse”) is detected. Progenitors followed by a vetoing pulse within 5 ms are discarded. We also require a progenitor to be at least 200 ms after any vetoing pulse to mitigate the aftermath of earlier energy depositions. Small pulses following these disqualified progenitors are not tracked in this analysis. Finally, events with otherwise valid progenitors are rejected if any data buffer for storing PMT digitizations is saturated. This cut effectively excludes progenitor pulses larger than phd and largely mitigates the effects of PMT saturation.
@@
After 3 ms, the pulses are almost exclusively SEs, the rate of which is more than an order of magnitude higher than that of pulses tagged as double electrons. Beyond a time delay of 10 ms, the rate of pulses identified as ME ( 3 electrons) are statistically consistent for the position-correlated and position-uncorrelated distributions, and do not exhibit significant time correlation with progenitors. This suggests that the dominant source of the ME background is independent of preceding energy depositions. The position-correlated SE rate, shown in Fig. 2, greatly exceeds that of position-uncorrelated SEs, so we attribute the position-correlated SEs to delayed emission related to earlier progenitor events. Similar to prior studies, the position-correlated SE rate approximately follows a power-law until contributions from position-uncorrelated pulses become significant. Therefore, we subtract the rate of position-uncorrelated pulses from the rate of position-correlated pulses in each bin, and fit the resulting distribution with the function , where and are constants, for times between 3 and 200 ms. At the nominal WS2022 field values we obtain a value of with negligible changes in value from varying the choice of fit bounds between 3–200 ms. When analyzing datasets acquired under the same nominal conditions but at times several months apart, we observe a 10% variation in , which we attribute to subtle changes in detector conditions and treat as a systematic error for this analysis. The obtained value and uncertainty level are in agreement with the values reported in LUX [27] and XENON1T [29], and will be discussed further in the next subsections.
@@ III.1 Impurity and drift time dependence
| Factor [# of Electrons] | Description |
| = | Extracted. "Raw" S2 size divided by SE size. |
| = | At liquid surface. divided by electron extraction efficiency, as calculated from [46] with a linear interpolation. |
| = | At interaction site. divided by drift survival probability. |
| = | Lost along drift path. |
@@
This explanation of the position-correlated delayed electron background has also been supported by experimental studies using LUX [27] and XENON1T [29] data. Because drift electrons are randomly captured by impurities along the drift path, the delayed release of captured electrons should occur independently, producing only delayed SEs. In this study, we observed S2 pulses with areas in the double electron (1.5–2.5 SE) region that follow large progenitor S2s, but their dependence on progenitor characteristics and detector conditions is nearly identical to that of SEs discussed above, and their rate is observed as a fixed percentage of the position-correlated SE rate. The same behaviors were also reported from the XENON1T experiment [29]. Therefore, we attribute these delayed “double-electron” backgrounds to either an upper fluctuation of the SE pulse area or the merge of an SE with its induced electron pulse. Examples of the latter effect include the photoionization of impurities near the liquid surface or photoelectric effects on the gate grids by electroluminescence photons from the initiating SE. Unlike XENON1T, position-uncorrelated electron backgrounds in this study are not observed to correlate with progenitor events, possibly a result of the accurate position reconstruction in LZ and our conservative cut choice for position-uncorrelated events.
@@ III.2 Electric field dependence
Table 2 lists all prior reports of the electron train power law in dual-phase TPCs. An exponent value ranging from -1 to -1.4 is reported across different experiments. Consistent with our analysis shown in Fig. 3, the liquid xenon purity, measured as the electron lifetime, does not seem to have significant bearing on the power law slope. In addition, these measurements were carried out with varying electron extraction efficiency values, which also do not seem to impact the power-law exponent. In LZ, we analyzed data acquired at a fixed drift field (193 V/cm) and three different extraction fields: 3410, 3900, and 4300 V/cm. These extraction fields correspond to extraction efficiencies of 50–78%. The power-law exponent values measured at these different field settings were consistent with each other.
@@
Due to the short drift distance for progenitor ionization signals, very few electrons are captured by impurities in this volume, resulting in low statistics of delayed electrons. However, after subtracting the position-uncorrelated electron rate, we observe clear power-law features up to a 20 ms delay time window. We fit power laws to the background-subtracted rates at both field settings with a fit range restricted to 3–20 ms. At 3.4 kV/cm, we obtain and at 3.9 kV/cm we obtain . These results, along with the others reported in this work, are summarized in Tab. 2.
@@
Unfortunately, there has been no compelling hypothesis for the mechanism behind delayed electron emission. The non-integer power-law exponent observed in LZ and other experiments makes it difficult to attribute this phenomenon to processes such as diffusion [48] or recombination [49], which can lead to or time dependence. Non-integer power laws are nevertheless observed in systems with energy traps of varying depths [50]. Taking the example of pumping out a vacuum chamber after it is exposed to air [51, 52], when there is only one residual gas species with a specific surface adsorption energy, the chamber pressure will experience an exponential decay when it is pumped on. However, when a range of adsorption energy values are introduced, the overall gas pressure exhibits a power-law time dependence. In liquid xenon, Ref. [34] proposes that thermal collisions with neighboring atoms may cause negatively charged impurities to be re-ionized and electrons to be released. In principle, these interactions should also result in an exponential time dependence for the electron emission rate, but the presence of multiple species of impurities with varying electron affinity can lead to an overall emission rate in the form of a power-law of non-integer exponents. Furthermore, the thermal energy of charged particles in an electric field depends on the field strength, so the hypothesis that charged impurities release electrons as a result of thermal collision can be tested by a definitive field-dependence study in the future.
@@ IV.1 Photons accompanying grid electron emission
This section studies electron emission from the LZ gate electrode using an extraction region-only dataset acquired during commissioning. This configuration greatly reduces delayed electron backgrounds following particle interactions in the drift region and is therefore ideal for characterizing grid emission. To further reduce the delayed electron background from interactions in the above-gate liquid, we veto time periods of following large S2s, where the veto length increases with the size of the S2 [3]. We also require that the S2 pulses do not have S1s larger than 100 phd within 3 s beforehand to exclude photoionization electrons. After applying this selection and additional position reconstruction quality cuts, the X-Y position distribution of the remaining isolated electron pulses is shown in Fig. 6 (top). Localization of the pulses around specific “hot spots,” or spontaneous emitters, is clearly visible. In addition to the spatial inhomogeneity, large temporal spikes in the S2 rates are also present in this data set, as shown in Fig. 6 (bottom). Because the cathode is inactive, the hot spots are attributed to electron emission from defects or debris on the gate electrode. During the enhanced electron emission periods, high rates of photons are also observed, a phenomenon also reported in LUX [37]. Figure 7 (top) shows the arrival time of photons relative to that of S2s, or the rate of single photo-electrons (SPEs) per S2 per second, as a function of the time difference between the SPE and S2. SPEs are defined in the LZ reconstruction algorithm as S1-like signals with light in only one PMT. A prominent peak is observed at 2.2 s, which is consistent with the drift time from the gate wire to the liquid surface at the WS2022 field. As described in Sec. III.2, the maximum drift time for extraction region events is approximately 2.5 s, but the drift time for gate-emitted electrons also depends on the exact emission position on the wire and the local field structure. This observation confirms that these photons and electrons are emitted around the same time and that the electrons originate from the gate grid. Similar gate photon-electron coincidences are also observed in a study of radiogenic backgrounds on LZ grids [39]; the rate of grid radioactive decays, however, is over an order of magnitude lower than that observed for field-induced emission in the few-electron region and thus its contribution to this analysis is neglected.
@@ IV.2 Implications for dark matter searches
The vast majority of grid-emitted S2 pulses observed in LZ (Fig. 8) contain fewer than 5 electrons (300 phd). The standard WIMP search of LZ uses an S2 threshold of approximately 10 electrons [3], so grid emission only contaminates the region of interest when rates are high enough to pile up significantly. In WS2022 [3], we implemented a rate-based hot spot exclusion cut by removing time intervals (120 seconds) where the average rate of low-energy S2s (including SEs) exceeded a certain threshold. This cut was demonstrated to be sufficient in reducing spurious S2 backgrounds for a sensitive search of medium to high-mass dark matter candidates.
@@
For low-mass dark matter searches, we usually focus on events with no identified S1s111The S1 definition is the same as that used in WS2022 where at least 3 photomultiplier tubes must observe light. within the 4.5 ms event window, and this cut is found to accept of S2-only events between 10 and 50 electrons. The signal loss is dominated by false SPE tagging, which we further investigate with a different method. The WS2022 data, with the TPC drift region active, contains a higher SPE rate than that in the extraction region-only dataset shown in Fig. 7. As a result, the probability for a SPE to be randomly paired with a S2 pulse can be significant and cause non-grid S2 pulses to be falsely rejected. We carried out a mock analysis by tagging SPEs arriving in a control drift time window that is equal in size but offset from the gate cut interval. Then we remove all S2 pulses in an event if any S2 pulse of 300 phd has a tagged SPE in the control coincidence window. The SPE-S2 pairs tagged in this way are only from random coincidence and the resulting event rate reduction is a measure of signal loss from random photon coincidence. When this mock cut is applied to S2-only events, we found a survival fraction of for events with S2 pulses above 300 phd (and below 2,880 phd, or 50 electrons), and a lower fraction of for events with S2 pulses below 300 phd. The survival fractions are different for S2 pulses above and below 300 phd because a large S2 can only be rejected in this cut if it shares the event window with a tagged S2 of 300 phd, but a small S2 can be additionally rejected if it has a preceding SPE pulse. Above 300 phd, there is minimal dependence on pulse area; the survival fraction for S2-only events above 10 electrons is .
The uncertainties quoted above are statistical only, and a full acceptance analysis is not attempted here because it would depend on data trigger efficiency and complex timing structures of SPE and S2 rates. The obtained efficiency values, nevertheless, suggest that a substantial reduction in grid backgrounds can be achieved with a high signal acceptance of 90% or more. It is worth noting that this proof-of-principle analysis does not use the standard LZ hot spot cut that excludes events recorded during periods of high average S2 rates. There is considerable overlap between the background populations removed by these two cuts. However, the event-level method developed in this work can enhance sensitivity to weak grid-emission with an average rate too low to be detected by rate-based cuts. This cut can be further optimized to balance background rejection power and signal loss by tuning the width of the pulse rejection window and the minimum number of photon-tagged S2s within the window. One can further exploit the position distribution of tagged S2s to further reduce accidental coincidences and to improve the signal efficiency. The optimization of this background rejection method can be explored in future work.
@@ Acknowledgements.
- Akerib et al. [2022] D. S. Akerib, P. B. Cushman, C. E. Dahl, R. Ebadi, A. Fan, R. J. Gaitskell, C. Galbiati, G. K. Giovanetti, G. B. Gelmini, L. Grandi, S. J. Haselschwardt, C. M. Jackson, R. F. Lang, B. Loer, D. Loomba, M. C. Marshall, A. F. Mills, C. A. J. OHare, C. Savarese, J. Schueler, M. Szydagis, V. Takhistov, T. M. P. Tait, Y. D. Tsai, S. E. Vahsen, R. L. Walsworth, and S. Westerdale, Snowmass2021 Cosmic Frontier Dark Matter Direct Detection to the Neutrino Fog (2022), arXiv:2203.08084 [hep-ex] .
@@
- Essig et al. [2023] R. Essig et al., Snowmass2021 Cosmic Frontier: The landscape of low-threshold dark matter direct detection in the next decade (2023), arXiv:2203.08297 [hep-ph] .
@@
- Li et al. [2023] S. Li et al. (PandaX Collaboration), Physical Review Letters 130, 10.1103/physrevlett.130.261001 (2023).
- Aprile et al. [2025] E. Aprile et al. (XENON Collaboration), Phys. Rev. Lett. 134, 161004 (2025).
- Aprile et al. [2026] E. Aprile et al. (XENON Collaboration), Light Dark Matter Search with 7.8 Tonne-Year of Ionization-Only Data in XENONnT (2026), arXiv:2601.11296 [hep-ex] .
@@
- Ibe et al. [2018] M. Ibe, W. Nakano, Y. Shoji, and K. Suzuki, JHEP 03, 194, arXiv:1707.07258 [hep-ph] .
@@
- Aprile et al. [2022a] E. Aprile et al. (XENON Collaboration), The European Physical Journal C 82, 10.1140/epjc/s10052-022-10345-6 (2022a).
@@
- Kopec et al. [2021] A. Kopec, A. Baxter, M. Clark, R. Lang, S. Li, J. Qin, and R. Singh, Journal of Instrumentation 16 (07), P07014.
- Aprile et al. [2022b] E. Aprile et al. (XENON Collaboration), Physical Review D 106, 10.1103/physrevd.106.022001 (2022b).
- Akimov et al. [2016] D. Akimov, V. Belov, A. Bolozdynya, A. Burenkov, Y. Efremenko, A. Etenko, V. Kaplin, A. Karelin, A. Khromov, M. Kirsanov, S. Klimanov, A. Kobyakin, A. Konovalov, A. Kovalenko, A. Kuchenkov, A. Kumpan, Y. Melikyan, R. Nikolaev, D. Rudik, G. Simakov, V. Sosnovtsev, and V. Stekhanov, Journal of Instrumentation 11 (03), C03007.
@@
- Sorensen and Kamdin [2018] P. Sorensen and K. Kamdin, Journal of Instrumentation 13 (02), P02032, arXiv:1711.07025 [astro-ph, physics:hep-ex, physics:physics].
@@
- Akerib et al. [2026] D. S. Akerib et al. (LZ Collaboration), Low-energy radon backgrounds from electrode grids in dual-phase xenon tpcs (2026), arXiv:2602.21177 [hep-ex] .
@@
- Mount et al. [2017] B. J. Mount et al. (LZ Collaboration), LUX-ZEPLIN (LZ) Technical Design Report (2017), arXiv:1703.09144 [physics.ins-det] .
@@
- [46] J. Xu, S. Pereverzev, B. Lenardo, J. Kingston, D. Naim, A. Bernstein, K. Kazkaz, and M. Tripathi, Phys. Rev. D 99, 103024.
@@
- Chiggiato [2020] P. Chiggiato, Outgassing properties of vacuum materials for particle accelerators (2020), arXiv:2006.07124 [physics.acc-ph] .